# Your copula is a classifier in disguise: classification-based copula density estimation

### v1

@@ 7 Discussion
Future directions of our work can incorporate different losses (Menon and Ong,, 2016) to fine-tune ratio copulas to focus on traits such as extreme dependence in the tails. An exciting direction is utilising other DRE methods for copula estimation, most notably linking copula density estimation with (time) score matching as in Choi et al., (2022). Crucially, given the success copulas have enjoyed in related fields (Tagasovska et al.,, 2019; Czado et al.,, 2022; Huk et al.,, 2024) as methodological tools, further improvements would yield dividends across all of their applications.
@@
David Huk originated the idea, designed the project, wrote the code, ran the experiments, wrote the paper and derived the proofs. Ritabrata Dutta and Mark Steel jointly supervised the project, provided valuable feedback, and helped edit the paper.
@@ Appendix A Proofs
We begin by stating assumptions needed for each part of the proof, followed by proving each of the three statements, namely [1] uniqueness, [2] consistency in the limit of infinite data from $p$, [3] asymptotic normality of the parameter estimate. This result is largely an application of Gutmann and Hyvärinen, (2012) Theorems 1,2 and 3.
@@
We begin by assuming that all marginals are known and, that they are continuous on the space $\mathcal{X}\in\mathbb{R}$ where $p(\mathbf{x})>0$ so that the copula exists by Sklar’s theorem Sklar, (1959). We consider the loss $\mathcal{L}(\theta)$ from (A.1), as in Gutmann and Hyvärinen, (2012).
@@
This is a classic assumption for deriving statistical results in model estimation, as done in the proof for the rate of convergence for simplified vine copula models Nagler and Czado, (2016). In our case, using MLP-based copula models gives us access to a much larger model class compared to simplified vines, making the discrepancy between the true model and the best model within our model class smaller, thereby reducing misspecification concerns.
@@
As our marginals are positive whenever $p(\mathbf{x})>0$, the same goes for $q(\mathbf{x})=\prod_{i=1}^{d}p_{i}(x_{i})$. This then applies to $\tilde{p}$ and $\tilde{q}$ too. Therefore, we satisfy the assumptions of Theorem 1 in Gutmann and Hyvärinen, (2012), which gives us the result namely that ${c}^{\text{r}atio}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x});\boldsymbol{\theta}^{*})$ is the only minimised of our loss. Further, as the model is well specified due to A2, we recover $c(\mathbf{x})=\tilde{p}(\mathbf{x})/\tilde{q}(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$.∎
@@
Keeping $\nu$ fixed, and increasing $T_{p}$ (meaning $T_{q}$ will also increase at a fixed constant rate), the supremum of the empirical approximation ^$\widehat{\mathcal{L}}(\boldsymbol{\theta})$ of Loss A.1 is such that, with the empirical loss defined as:
@@
Assumptions A3 and A4 have their counterparts in maximum likelihood estimation, see e.g. Wasserman, (2006).
@@
As assumptions A1, A2, A3 and A4 imply that we satisfy assumption (a,b,c) of theorem 2 in Gutmann and Hyvärinen, (2012), we obtain its result, namely that $\widehat{\boldsymbol{\theta}}\rightarrow^{\mathbb{P}}\boldsymbol{\theta}^{*}$. ∎
@@
As we satisfy assumption A1-4, theorem 3 in Gutmann and Hyvärinen, (2012) applies, and we obtain that $\sqrt{T_{p}}\left(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}^{\star}\right)$ is asymptotically normal with mean $\mathbf{0}$ and covariance matrix $\boldsymbol{\Sigma}$ wit expression:
@@
In particular, as $\nu\rightarrow\infty$, $\Sigma$ becomes independent of the form of $\tilde{q}$, meaning that no matter the complexity of $c(\cdot)$ relative to a simple independence copula, with infinite noise samples, estimating it is only dependent on the shape of $p$ alone. Furthermore, under some mild assumptions detailed in Geenens et al., (2017), the resulting $\tilde{p}$ after applying transformation (5) has unconstrained support and it, as well as its partial derivatives, will be bounded on $\mathbb{R}^{d}$. This is useful to ensure a well-behaved $\Sigma$ for asymptotic normality, since $\Sigma$ is defined in terms of the gradients of $\tilde{p}$.
@@
A.4 Proposition 4.2
@@
A.5 Corollary 4.3
We begin by stating the simplified vine copula density, in broad terms. In the equation below, we use $u_{i}$ and $u_{j}$ as copula observations from two distinct dimensions $i,j\in\{1,\ldots,d\}$, possibly involving conditioning on other variables aside from $i,j$. For a more detailed version of the likelihood, the reader is referred to Equation 29 in Appendix B.
$\mathbf{c}(u_{1},\ldots,u_{d})=\prod_{i\neq j}^{d(d-1)/2}c_{ij}({u_{i}},{u_{j}})$ (24)
Each copula takes two arguments $u_{i},u_{j}\in[0,1]$as inputs, and outputs their bivariate copula density. Based on the model chosen for $c_{i,j}$, we can deduce separate equivalences for the final vine model.
@@
When using bivariate copula building blocks, the vine equation becomes
$\mathbf{c}(u_{1},\ldots,u_{d})=\prod_{i\neq j}^{d(d-1)/2}\frac{\mathcal{N}_{2}(\Phi^{-1}(u_{i}),\Phi^{-1}(u_{j});\mathbf{0},\mathbf{\Sigma})}{\mathcal{N}(\Phi^{-1}(u_{i});{0},{1})\cdot\mathcal{N}(\Phi^{-1}(u_{j});{0},{1})}.$ (25)
This expression is a ratio of two bivariate Gaussians with means $\mathbf{0}$ and covariances $\mathbf{\Sigma},\mathbf{I}_{2}$ for the numerator and denominator respectively. From Proposition 4.1, it follows that each term in the vine decomposition is a QDA decision boundary, giving the result.
@@
Similarly, each bivariate KDE copula is unconditional, with some inputs on the $\mathbb{R}^{2}$ Gaussian latent space. This implies a ratio, and so Proposition 4.2 applies, meaning that the simplified vine density is then composed as a product of $\frac{d(d-1)}{2}$ bivariate classifiers, each being an average of $T_{p}$ decision boundaries obtained by QDA on a 2-dimensional Gaussian problem.
@@ B.2 Vine copulas
A vine copula is a flexible method for modelling dependencies in high-dimensional data by decomposing a $d$-dimensional copula into a sequence of $d(d-1)/2$ bivariate copulas through structured conditioning (Bedford and Cooke,, 2001). This allows for efficient estimation by modelling pairs of variables instead of the full joint distribution (Nagler and Czado,, 2016). The decomposition is based on the following identity, which follows from Sklar’s theorem:
@@ Appendix C Practical details
    • Calculate loss $\mathcal{L}(\boldsymbol{\theta})$ using Equation (A.1);
@@
In our experiments, for the simplified vine copula, we use a TLL constant copula as the bivariate building block, corresponding to a bivariate KDE copula. The TLL constant copula requires a bandwidth $b$ for the kernel. We followed standard literature on KDE estimation by choosing Silverman’s rule of thumb for the form of the bandwidth (Silverman,, 2018), based on the sample size $n$ in the training data, and the dimension of the data $d$:
@@
However, in the Digits experiment, this choice resulted in numeric overflow in the LL evaluation (likely due to the low number of sampled compared to the dimension), leading us to choose the bandwidth $b$ with a grid search using cross-validation on training data. Further, simplified vine copulas can be truncated at a given level $t\leq d$, meaning that the number of bivariate copulas is limited at the first $t(t-1)/2$ such copulas in the decomposition. This is equivalent to setting all bivariate copulas past a certain level of conditioning as independent copulas (i.e equal to $1$). This is a common approach to reduce the computational burden of fitting a vine copula (Czado and Nagler,, 2022). We use a full set of $d(d-1)/2$ bivariate copulas for the Digits experiment, but use a truncation level of $50$ for the MNIST experiment, in order to keep computational times of the vine copula similar to other models.
For the IGC, we follow the experimental setup of Janke et al., (2021) in their original paper. We use a fully connected MLP, with an initial layer taking as input Gaussian noise, with dimension $3d$, followed by two fully connected layers with width $100$, with a final connection from the second fully connected layer to the $d$-dimensional output. This architecture is used for all experiments with the exception of MNIST where a third fully connected hidden layer was used. We show the parameter number per experiment in Table 2.
@@
Finally, our ratio copula model follows a similarly simple architecture to the IGC, to show its general applicability. We specifically chose architectures with a similar parameter count to make comparisons fair. In the Digits and MNIST examples, we use convolutional layers following a standard classifier architecture for image data. We provide a table describing our ratio copula architecture in Tables 3, 4, 5 and 6. We use a different number of independent noise samples in the optimisation procedure of Loss (A.1) based on the complexity of the task and the desired trade-off between computational time and statistical efficiency. In the 2D parametric copula example, we use the whole 5000 samples and 50,000 noise samples (but these are fixed across epochs). In the 2D image experiment, we use 10,000 true samples and 100,000 noise samples per epoch. In the Digits experiment, we use the whole 898 training samples and use 89,800 noise samples per batch. In the MNIST example, we use the whole data of size 30,000 and 300,000 independent samples.
@@
For the IGC and ratio copula models, we employ the Adam optimiser (Diederik,, 2014) in Pytorch with the base settings, to demonstrate the ease of use of our model, as in Janke et al., (2021). For the 2D parametric copula experiments, we optimise these models for 500 steps, for the 2D image experiments we use 100,000 optimisation steps, for the Digits dataset we use 500 steps and we use 1000 steps for MNIST to reflect the complexity of the task. For the ratio copula, at each epoch, we resample a new set of independent noise (either from $\mathcal{N}(\cdot;\mathbf{0},\mathbf{I}_{d})$ or from a correlated Gaussian $\mathcal{N}(\cdot;\mathbf{0},\Sigma)$ in the Gaussian guided ratio copula). Therefore, for the ratio copula, following the analysis in Theorem 3.1 and Gutmann and Hyvärinen, (2012), there is a trade-off available for our method between computation time and statistical performance; as we optimise the model for more epochs and sample more independent noise, the better our estimate of the true ratio copula. Therefore, while we limit the number of epochs to the same number as the IGC, our method could benefit from longer training times.
@@
As our ratio copula model has an analytical density, we employ Markov chain Monte Carlo methods to obtain samples. Specifically, we use the Hamiltonian Monte Carlo scheme (Neal,, 2011), from the implementation in the pyhmc python package (McGibbon,, 2015). For each sample, we initiate the chain at a random sample from a standard independent Gaussian on the latent Gaussian space, and implement 100 gradient-informed MCMC steps to ensure the chain has reached the main mass of the density. We then select the next step after that as the sample from our model. As in e.g. KDE or Gaussian copula models, this sampling is done from $p$ on the latent Gaussian space, meaning that we multiply our final ratio copula by $\mathcal{N}(\cdot;\mathbf{0},\mathbf{I}_{d})$, and target the resulting density with our MCMC sampling scheme. The samples obtained by the MCMC scheme are consequently on a Gaussian scale, meaning we apply a Gaussian CDF to map them to the copula $[0,1]^{d}$ space.
@@
We use the LL, as measured by the output of our ratio model’s log-density on held-out test data. Similarly, we use the python implementation of Vatter and Nagler, (2022) of simplified vine copula models as well as bivariate copulas to output LL values. For the Wasserstein 2 metric, we make use of the implementation of Flamary et al., (2021). For all experiments, we measure the W2 using 500 samples from models, except for MNIST where we use 25 samples per model.

### v2

@@ 7 DISCUSSION
Future directions of our work can incorporate different losses [Menon and Ong, 2016] to fine-tune ratio copulas to focus on traits such as extreme dependence in the tails. An exciting direction is utilising other DRE methods for copula estimation, most notably linking copula density estimation with (time) score matching as in Choi et al. [2022]. Furthermore, some applications of copulas such as Bayesian Copula Networks Elidan [2010], and conformal prediction [Messoudi et al., 2021, Park et al., 2024] desire to work with the copula distribution rather than the density. Extending our ratio copula formulation to obtain a model that parameterises CDFs is another meaningful direction.
Crucially, given the success copulas have enjoyed in related fields [Tagasovska et al., 2019, Czado et al., 2022, Huk et al., 2024] as methodological tools, further improvements would yield dividends across all of their applications.
@@
David Huk originated the idea, designed the project, wrote the code, ran the experiments, wrote the paper, and derived the proofs. Ritabrata Dutta and Mark Steel jointly supervised the project, provided valuable feedback, and helped edit the paper.
Acknowledgements:
We thank all the reviewers for their helpful feedback. We also thank Alexis Derumigny, and Lorenzo Pacchiardi for helpful comments before the camera-ready version. David Huk is funded by the Center for Doctoral Training in Mathematical Sciences at Warwick. Ritabrata Dutta is funded by EPSRC (grant nos. EP/V025899/1 and EP/T017112/1) and NERC (grant no. NE/T00973X/1).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] We include a notation part in Section 2, and state assumptions at the start of relevant sections of the text.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] We provide Lemma 3.2 to quantify the complexity of our task.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] Will be included in the additional material and published upon acceptance.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes] In the Appendix.
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A PROOFS
We begin by stating assumptions needed for each part of the proof, followed by proving each of the three statements, namely [1] uniqueness, [2] consistency in the limit of infinite data from $p$, [3] asymptotic normality of the parameter estimate. This result is largely an application of Gutmann and Hyvärinen [2012] Theorems 1,2 and 3.
@@
We begin by assuming that all marginals are known and, that they are continuous on the space $\mathcal{X}\in\mathbb{R}$ where $p(\mathbf{x})>0$ so that the copula exists by Sklar’s theorem [Sklar, 1959]. We consider the loss $\mathcal{L}(\theta)$ from (3.1), as in Gutmann and Hyvärinen [2012].
@@
This is a classic assumption for deriving statistical results in model estimation, as done in the proof for the rate of convergence for simplified vine copula models Nagler and Czado [2016]. In our case, using MLP-based copula models gives us access to a much larger model class compared to simplified vines, making the discrepancy between the true model and the best model within our model class smaller, thereby reducing misspecification concerns.
@@
As our marginals are positive whenever $p(\mathbf{x})>0$, the same goes for $q(\mathbf{x})=\prod_{i=1}^{d}p_{i}(x_{i})$. This then applies to $\tilde{p}$ and $\tilde{q}$ too. Therefore, we satisfy the assumptions of Theorem 1 in Gutmann and Hyvärinen [2012], which gives us the result namely that ${c}^{\text{r}atio}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x});\boldsymbol{\theta}^{*})$ is the only minimised of our loss. Further, as the model is well specified due to A2, we recover $c(\mathbf{x})=\tilde{p}(\mathbf{x})/\tilde{q}(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$.∎
@@
Keeping $\nu$ fixed, and increasing $T_{p}$ (meaning $T_{q}$ will also increase at a fixed constant rate), the supremum of the empirical approximation ^$\widehat{\mathcal{L}}(\boldsymbol{\theta})$ of Loss 3.1 is such that, with the empirical loss defined as:
@@
Assumptions A3 and A4 have their counterparts in maximum likelihood estimation, see e.g. Wasserman [2006].
@@
As assumptions A1, A2, A3 and A4 imply that we satisfy assumption (a,b,c) of theorem 2 in Gutmann and Hyvärinen [2012], we obtain its result, namely that $\widehat{\boldsymbol{\theta}}\rightarrow^{\mathbb{P}}\boldsymbol{\theta}^{*}$. ∎
@@
As we satisfy assumption A1-4, theorem 3 in Gutmann and Hyvärinen [2012] applies, and we obtain that $\sqrt{T_{p}}\left(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}^{\star}\right)$ is asymptotically normal with mean $\mathbf{0}$ and covariance matrix $\boldsymbol{\Sigma}$ wit expression:
@@
In particular, as $\nu\rightarrow\infty$, $\Sigma$ becomes independent of the form of $\tilde{q}$, meaning that no matter the complexity of $c(\cdot)$ relative to a simple independence copula, with infinite noise samples, estimating it is only dependent on the shape of $p$ alone. Furthermore, under some mild assumptions detailed in Geenens et al. [2017], the resulting $\tilde{p}$ after applying transformation (5) has unconstrained support and it, as well as its partial derivatives, will be bounded on $\mathbb{R}^{d}$. This is useful to ensure a well-behaved $\Sigma$ for asymptotic normality, since $\Sigma$ is defined in terms of the gradients of $\tilde{p}$.
@@
A.4 Student’s t copula
We provide here a result on non-Gaussian copulas, where we use the definition of a student’s t-distribution with multiple degrees of freedom from Ogasawara [2024] (note that this is not the ”usual” multivariate Student with one degree of freedom parameter, where the common latent mixing variable induces dependence even if correlations are zero):
Proposition A.1.
Consider a Student’s t copula with density
$c_{\nu,P}^{t}(\mathbf{u})=\frac{f_{\nu,P}\left(F_{\nu}^{-1}\left(P_{1}(x_{1})\right),\ldots,F_{\nu}^{-1}\left(P_{d}(x_{d})\right)\right)}{\prod_{i=1}^{d}f_{\nu}\left(F_{\nu}^{-1}\left(P_{i}(x_{i})\right)\right)},$
where for $\nu$ the degrees of freedom, $f_{\nu,P}$ is the d-dimensional student’s t pdf with mean $\mathbf{0}$ and correlation matrix $P$, $F_{\nu}^{-1}$ is the inverse of a univariate standard student’s t cdf, and $f_{\nu}$ is the density of the univariate standard $t$-distribution. The associated Bayes optimal classifier of this copula is found by the decision boundary between a student’s t distribution and a student’s t distribution with multiple degrees of freedom when all its degrees of freedom are equal to $\nu$, both densities with mean $\mathbf{0}$ but correlation matrices $P$ and $diag(P)$ respectively.
Proof.
The decision boundary of the above-mentioned classification problem will take the form of a ratio between the two densities. The numerator is simply $f_{\nu,P}$, matching the numerator of the copula density. The denominator of the copula is a product of independent marginal student-t densities with common degrees of freedom $\nu$. We note that a multivariate student’s t does not factorise into $d$ marginals when its correlation matrix is diagonal. This can be understood through the equivalent representation of a student’s t with $\mathbf{Z}\sim f_{\nu,P}\iff\mathbf{Z}=\mathbf{X}/\sqrt{Y/\nu}$ where $\mathbf{X}\sim\mathcal{N}(\mathbf{0},P)$ and $Y$ follows a Chi-squared distribution with $\nu$ degrees of freedom. As the same $Y$ is used across all dimensions, the resulting $\mathbf{Z}$ variables are not marginally independent. However, this is the case in a multivariate t-distribution with multiple degrees of freedom with a diagonal correlation matrix, as different Chi-squared variables are used for each dimension. When all Chi-squared variables share the same degree of freedom, the factorised form becomes that of the denominator of a student’s t copula, concluding our proof. ∎
A.5 Proposition 4.2
@@ A.4 Student’s t copula
A.6 Corollary 4.3
Proof.
We begin by stating the simplified vine copula density for a $d$-dimensional vector $(u_{1},\ldots,u_{d})\in[0,1]^{d}$ and $i,j\in\{1,\ldots,d\}$. For more details on vine copulas, the reader is referred to Appendix B.
The conditioning sets $\mathcal{S}_{ij}$ are determined by the specific vine structure, and are a possibly empty subset of dimensions excluding $i,j$. Each of the $d(d-1)/2$ copulas in the decomposition takes two arguments $P_{i|\mathcal{S}_{ij}}(x_{i|\mathcal{S}_{ij}}),P_{j|\mathcal{S}_{ij}}(x_{j|\mathcal{S}_{ij}})$ as inputs, both marginally uniform in $[0,1]$, and outputs their bivariate copula density. Based on the modelling assumption chosen for those bivariate densities $c_{i,j}$, we follow our previous results to deduce separate equivalences for the final vine model.
@@
When using bivariate copula building blocks, the vine equation becomes:
$\mathbf{c}(u_{1},\ldots,u_{d})=\prod_{i\neq j}^{d(d-1)/2}\frac{\mathcal{N}_{2}\left(\Phi^{-1}\left(P_{i|\mathcal{S}_{ij}}(x_{i|\mathcal{S}_{ij}})\right),\Phi^{-1}\left(P_{j|\mathcal{S}_{ij}}(x_{j|\mathcal{S}_{ij}})\right);\mathbf{0},\mathbf{\Sigma}\right)}{\mathcal{N}(\Phi^{-1}\left(P_{i|\mathcal{S}_{ij}}(x_{i|\mathcal{S}_{ij}})\right);{0},{1})\cdot\mathcal{N}(\Phi^{-1}\left(P_{j|\mathcal{S}_{ij}}(x_{j|\mathcal{S}_{ij}})\right);{0},{1})}.$ (26)
As $P_{i|\mathcal{S}_{ij}}(x_{i|\mathcal{S}_{ij}}),P_{j|\mathcal{S}_{ij}}(x_{j|\mathcal{S}_{ij}})$ are marginally uniform, by the probability integral transformation, this expression is a ratio of two bivariate Gaussians with means $\mathbf{0}$ and covariances $\mathbf{\Sigma},\mathbf{I}_{2}$ for the numerator and denominator respectively. From Proposition 4.1, it follows that each term in the vine decomposition is a QDA decision boundary, giving the result.
@@
Similarly, each bivariate KDE copula is a ratio between a mixture of bivariate kernels and a bivariate standard Gaussian, with some inputs on the $\mathbb{R}^{2}$ Gaussian latent space that are marginally Gaussian due to the probability integral transformation. Due to the form of a Gaussian KDE copula, Proposition 4.2 applies to each of the $d(d-1)/2$ bivariate copulas involved in the vine decomposition (up to replacing the observations $\{(Z_{1}^{k},\ldots,Z_{d}^{k})\}_{k=1}^{T_{p}}$ by new ${Z_{1}^{\prime}}^{k}:=\Phi^{-1}\circ P_{i|\mathcal{S}_{ij}}(x_{i|\mathcal{S}_{ij}})$ which include the conditioning implied by vines), meaning that the simplified vine density is then composed as a product of $d(d-1)/2$ bivariate classifiers, each being an average of $T_{p}$ decision boundaries obtained by QDA on a 2-dimensional Gaussian problem. ∎
A.7 Equivalences for parametric copulas in terms of classifiers
We can use our framework to obtain the classifiers implied by known copulas. Indeed,
$c(\mathbf{P(\mathbf{x})})=\frac{h(\mathbf{x})}{1-h(\mathbf{x})}\iff h(\mathbf{x})=\frac{c(\mathbf{P(\mathbf{x})})}{1+c(\mathbf{P(\mathbf{x})})}$
which we use to empirically illustrate the classifiers implied by various copula families by using their densities.
Figure 4 shows six different bivariate copulas and their classifiers on the copula $[0,1]^{2}$ space as well as the Gaussian space via a dimension-wise inverse Gaussian transform. A notable benefit of representing copulas by their classifier counterparts is an improved clarity of the shape; the LL values $\log c(U,V)$ in the plots of all but the Frank copula explode at the corners which makes them hard to evaluate, whereas the classifier versions $h(u,v)$ depict clear shapes in accordance with those of copulas. This can be seen on the Gaussian space plots, where the classifier identifies the diagonal as being high-likelihood regions for the dependent data and places increasingly high probability on the diagonal as the distance from the centre (and so from the independent factorised density) increases. The classifier’s increasingly high probability is expected, as for instance, the Gaussian copula density is unbounded in the corners on the diagonal. Using our classifier interpretation also exemplifies the theoretical results from Section 4. Further, the similarity of the classifier to the shape of the copula is expected as the copula LL can be understood as a difference of the LL of the numerator and denominator of a copula i.e. $\log c(P_{1}(c_{1}),P_{2}(x_{2}))=\log p(x_{1},x_{1})-\log(p_{1}(x_{1})\cdot p_{2}(x_{2}))$, which is what the classifier is trained to distinguish.
Figure 4: Example of six parametric bivariate copulas and their equivalent classifiers $h^{*}(\mathbf{x})$. The first two columns show the copula log-likelihoods (LL) and their implied classifiers on the copula $[0,1]^{2}$ scale, while the last two columns show the same quantities but on the Gaussian $\mathbb{R}^{2}$ scale.
@@ B.2 Vine copulas
A vine copula is a flexible method for modelling dependencies in high-dimensional data by decomposing a $d$-dimensional copula into a sequence of $d(d-1)/2$ bivariate copulas through structured conditioning [Bedford and Cooke, 2001]. This allows for efficient estimation by modelling pairs of variables instead of the full joint distribution [Nagler and Czado, 2016]. The decomposition is based on the following identity, which follows from Sklar’s theorem:
@@ Appendix C PRACTICAL DETAILS
    • Calculate loss $\mathcal{L}(\boldsymbol{\theta})$ using Equation (3.1);
@@
In our experiments, for the simplified vine copula, we use a TLL constant copula as the bivariate building block, corresponding to a bivariate KDE copula. The TLL constant copula requires a bandwidth $b$ for the kernel. We followed standard literature on KDE estimation by choosing Silverman’s rule of thumb for the form of the bandwidth [Silverman, 2018], based on the sample size $n$ in the training data, and the dimension of the data $d$:
@@
However, in the Digits experiment, this choice resulted in numeric overflow in the LL evaluation (likely due to the low number of sampled compared to the dimension), leading us to choose the bandwidth $b$ with a grid search using cross-validation on training data. Further, simplified vine copulas can be truncated at a given level $t\leq d$, meaning that the number of bivariate copulas is limited at the first $t(t-1)/2$ such copulas in the decomposition. This is equivalent to setting all bivariate copulas past a certain level of conditioning as independent copulas (i.e equal to $1$). This is a common approach to reduce the computational burden of fitting a vine copula [Czado and Nagler, 2022]. We use a full set of $d(d-1)/2$ bivariate copulas for the Digits experiment, but use a truncation level of $50$ for the MNIST experiment, in order to keep computational times of the vine copula similar to other models.
For the IGC, we follow the experimental setup of Janke et al. [2021] in their original paper. We use a fully connected MLP, with an initial layer taking as input Gaussian noise, with dimension $3d$, followed by two fully connected layers with width $100$, with a final connection from the second fully connected layer to the $d$-dimensional output. This architecture is used for all experiments with the exception of MNIST where a third fully connected hidden layer was used. We show the parameter number per experiment in Table 2.
@@
Finally, our ratio copula model follows a similarly simple architecture to the IGC, to show its general applicability. We specifically chose architectures with a similar parameter count to make comparisons fair. In the Digits and MNIST examples, we use convolutional layers following a standard classifier architecture for image data. We provide a table describing our ratio copula architecture in Tables 3, 4, 5, 6, 7, and 8. We use a different number of independent noise samples in the optimisation procedure of Loss (3.1) based on the complexity of the task and the desired trade-off between computational time and statistical efficiency. In the 2D parametric copula example, we use the whole 5000 samples and 50,000 noise samples (but these are fixed across epochs). In the 2D image experiment, we use 10,000 true samples and 100,000 noise samples per epoch. In the Digits experiment, we use the whole 898 training samples and use 89,800 noise samples per batch. In the MNIST example, we use the whole data of size 30,000 and 300,000 independent samples.
@@
Linear [-1, 256] 65,792
LeakyReLU [-1, 256] 0
Linear [-1, 1] 257
Total params: 103,617
Table 6: Summary of the GG+Ratio large copula architecture used in Digits experiments.
Layer (type) Output Shape Param #
Conv2d [-1, 64, 14, 14] 640
LeakyReLU [-1, 64, 14, 14] 0
Conv2d [-1, 64, 7, 7] 36,928
LeakyReLU [-1, 64, 7, 7] 0
Flatten [-1, 3136] 0
@@
Layer (type) Output Shape Param #
Conv2d [-1, 64, 14, 14] 640
LeakyReLU [-1, 64, 14, 14] 0
Conv2d [-1, 64, 7, 7] 36,928
LeakyReLU [-1, 64, 7, 7] 0
Flatten [-1, 3136] 0
Linear [-1, 128] 401,536
LeakyReLU [-1, 128] 0
Linear [-1, 64] 8,256
LeakyReLU [-1, 64] 0
Linear [-1, 64] 4,160
Linear [-1, 32] 2,080
Linear [-1, 1] 33
Total params: 453,633
Table 8: Summary of the GG+ratio large model architecture used in MNIST experiments.
@@
For the IGC and ratio copula models, we employ the Adam optimiser [Diederik, 2014] in Pytorch with the base settings, to demonstrate the ease of use of our model, as in Janke et al. [2021]. For the 2D parametric copula experiments, we optimise these models for 500 steps, for the 2D image experiments we use 100,000 optimisation steps, for the Digits dataset we use 500 steps and we use 1000 steps for MNIST to reflect the complexity of the task. For the ratio copula, at each epoch, we resample a new set of independent noise (either from $\mathcal{N}(\cdot;\mathbf{0},\mathbf{I}_{d})$ or from a correlated Gaussian $\mathcal{N}(\cdot;\mathbf{0},\Sigma)$ in the Gaussian guided ratio copula). Therefore, for the ratio copula, following the analysis in Theorem 3.1 and Gutmann and Hyvärinen [2012], there is a trade-off available for our method between computation time and statistical performance; as we optimise the model for more epochs and sample more independent noise, the better our estimate of the true ratio copula. Therefore, while we limit the number of epochs to the same number as the IGC, our method could benefit from longer training times.
Diagnostics for overfitting
Other than using Lemma 3.2, an easily available diagnostic is to look at the loss during training. Since the copula naturally overlaps independent with dependent samples, we should always observe a positive loss. If it is too close to 0, then the classifier is overfitting and the two densities are too far apart [Rhodes et al., 2020]. Furthermore, using a normalising constant estimator also provides a diagnostic of fit: copula densities are already normalised by definition, hence if correctly estimated, the normalising constant will be 1. If any of these symptoms arise, it signals we should employ a better DRE method for estimating the ratio copula. For instance, switching from a ratio copula to a Gaussian-Guided ratio copula.
@@
As our ratio copula model has an analytical density, we employ Markov chain Monte Carlo methods to obtain samples. Specifically, we use the Hamiltonian Monte Carlo scheme [Neal, 2011], from the implementation in the pyhmc python package [McGibbon, 2015]. For each sample, we initiate the chain at a random sample from a standard independent Gaussian on the latent Gaussian space, and implement 100 gradient-informed MCMC steps to ensure the chain has reached the main mass of the density. We then select the next step after that as the sample from our model. As in e.g. KDE or Gaussian copula models, this sampling is done from $p$ on the latent Gaussian space, meaning that we multiply our final ratio copula by $\mathcal{N}(\cdot;\mathbf{0},\mathbf{I}_{d})$, and target the resulting density with our MCMC sampling scheme. The samples obtained by the MCMC scheme are consequently on a Gaussian scale, meaning we apply a Gaussian CDF to map them to the copula $[0,1]^{d}$ space.
@@
We use the LL, as measured by the output of our ratio model’s log-density on held-out test data. Similarly, we use the python implementation of Vatter and Nagler [2022] of simplified vine copula models as well as bivariate copulas to output LL values. For the Wasserstein 2 metric, we make use of the implementation of Flamary et al. [2021]. For all experiments, we measure the W2 using 500 samples from models, except for MNIST where we use 25 samples per model.